OpenStudy (anonymous):

Simplify 2 to 4 power + 3 to the second power

OpenStudy (anonymous):

so you add them all together

OpenStudy (anonymous):

so the answer is 14

OpenStudy (anonymous):

@juanpabloJR

OpenStudy (anonymous):

Answer is 19! @diannemarieee

OpenStudy (anonymous):

ok thank you soo much

OpenStudy (anonymous):

19 isnt a chocie

OpenStudy (anonymous):

the answer is 25

OpenStudy (anonymous):

Since to is to the power of 4 it means that 2^4= 2*2*2*2= 16 and when u add 3 u get 19

OpenStudy (anonymous):

*twp

OpenStudy (anonymous):

but its asking to simplify

OpenStudy (anonymous):

**two

OpenStudy (anonymous):

Oh sorry didn't read the whole question!

OpenStudy (anonymous):

Got it! its like this ' 2^4 + 3^2= 16 + 9= 25' :)